# Sample complexity of optimal transport barycenters with discrete support

### v2

@@ 4 Sample complexity for sparse optimal transport barycenters
In this section, we present our main results : the generalization error bound followed by the sample complexity of classical, entropic and sliced optimal transport barycenters. These were inspired by (gyorfi2002principles, Section 4.4), which provides the sample complexity of the K-means functional (see Section 5.4).
@@
where the finite constant values $C_{p,R}$ and $C_{d,N}$ depend on the divergence $\mathbb{D}$ and are summarized in the following Table 2.
@@
Theorem 4.1 provides uniform bounds for the empirical barycenter cost function, which is the cost function minimized by most algorithms in practice when computing an optimal transport barycenter (See Section 3).
@@
Note that we consider here empirical measures supported on a set of samples of equal size $n$, but our results also hold for different sample sizes for each target measure by replacing $n$ by the minimum of those individual sample sizes in the upper bound of the generalization error. Finally, although for convenience we consider the barycenter problem (2.6) with uniform weights $\frac{1}{L}$ for each measure $\mu^{\ell}$, Theorem 4.1 holds for any weights $\lambda\in\Delta_{L}$.
A sketch of proof is presented below, with the complete details provided in Appendix C.
@@
We treat here the case $\mathbb{D}=W_{p}^{p}$ for simplicity. Let $(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}$. The first step is to control, for each $\ell=1,\ldots,L$, the deviation
$\left|W_{p}^{p}\!\left(\mu_{n}^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-W_{p}^{p}\!\left(\mu^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|.$
This can be achieved by resorting to the dual formulation of $W_{p}^{p}$, in which the dual variables are vectors $w\in\mathbb{R}^{N}$. Lemma D.2 then ensures that there exists functions $f_{Y}^{w}$ such that the above quantity can be upper bounded for each $\ell$ by
$\sup_{w\in\mathbb{R}^{N}}\left|\mathbb{E}_{X\sim\mu_{n}^{\ell}}\!\left[f_{Y}^{w}(X)\right]-\mathbb{E}_{X\sim\mu^{\ell}}\!\left[f_{Y}^{w}(X)\right]\right|.$
At first sight, one cannot apply classical tools from empirical process theory, since the dual variables $w$ are unbounded. Fortunately, Proposition B.1 guarantees that it is enough to restrict the supremum to $w\in[-K,K]^{N}$ for some constant $K$ independent of $Y$ and $\pi$. This boundedness is crucial, as it allows us to invoke standard results from empirical process theory (e.g., (gine2021mathematical, Theorem 3.5.1)) to conclude. ∎
@@
In the specific case where $\mathbb{D}$ is a distance, using the triangular inequality we have that
$|\mathbb{D}(\mu_{n}^{\ell},\nu)-\mathbb{D}(\mu^{\ell},\nu)|\leq\mathbb{D}(\mu_{n}^{\ell},\mu^{\ell}).$ (4.1)
Results on sample complexity of $\mathbb{D}$ could then be used to derive an upper bound on the estimation error in Theorem 4.1. However, for $\mathbb{D}=W_{1}$ for example, the right hand side term of (4.1) is of order $n^{-1/d}$ (chewi2024statistical, Proposition 2.1), which is worse than $n^{-1/2}$ when $d>1$.
Theorem 4.1 directly yields the sample complexity of sparse optimal transport barycenters, as stated in the following corollary.
@@
Let $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{B}_{R})$, for $R>0$. For some integer $n$, let $\mu_{n}^{1},\ldots,\mu_{n}^{L}$ be empirical measures supported over $n$ i.i.d random variables of respective law $\mu^{\ell}$, for all $\ell\in[\![1,L]\!]$. Let also $N\geq 1$ be some integer, $A$ a closed nonempty subset of $\mathbb{B}_{R}^{N}\times\Delta_{N}$, and let $\mathbb{D}=W_{p}^{p},\>W_{\epsilon,p}^{p}$ (for some $\epsilon>0$), $SW_{p}^{p}$ or $\text{max-}SW_{p}^{p}$. We denote by $(Y_{n},\pi_{n})$ an argmin in $A$ of
$(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=\frac{1}{L}\sum_{\ell=1}^{L}\mathbb{D}\left(\mu_{n}^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right).$
Then we have :
$\begin{split}&\mathbb{E}\left[\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}\right)-\underset{(Y,\pi)\in A}{\min}\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right]\\ &\leq 2C_{p,R}\sqrt{\frac{C_{d,N}}{n}},\end{split}$ (4.2)
@@
For the specific case of $\mathbb{D}=W_{2}^{2}$ and $L=1$, it was shown in biau2008performance that the estimation error is upper bounded by $O(N/\sqrt{n})$, but with no multiplicative dependence on the dimension $d$. However, this bound is worse by a factor $\sqrt{N}$ compared to our results, for which the rate is of order $O(\sqrt{dN/n})$. The proof of our Corollary 4.4 is obtained leveraging a classical statistical learning theory argument given in the proof below.
@@
Suppose $A$ is a closed subset of the compact set $\mathbb{B}_{R}^{N}\times\Delta_{N}$. By continuity of the functionals $(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}})$ and $(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}})$, they both admit global minimizers in $A$. We denote by $(Y^{*},\pi^{*})$ and $(Y_{n}^{*},\pi_{n}^{*})$ two respective minimizers, and their associated discrete measures $\nu^{*}:=\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}$ and $\nu_{n}^{*}:=\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}$. Then we have
@@
Theorem 4.1 and the sample complexity results of Sinkhorn divergences in genevay2019sample also yield an upper bound for the debiased Sinkhorn barycenters, that is for $\mathbb{D}=\overline{W}_{\epsilon,p}$ defined in (2.4). The proof of the next corollary is also postponed in Appendix C.
@@
Let $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{B}_{R})$, for $R>0$. For some integer $n$, let $\mu_{n}^{1},\ldots,\mu_{n}^{L}$ be empirical measures supported over $n$ i.i.d random variables of respective law $\mu^{\ell}$, for all $\ell\in[\![1,L]\!]$. Let also $N\geq 1$ be an integer, $A$ a closed nonempty subset of $\mathbb{B}_{R}^{N}\times\Delta_{N}$ and let $\mathbb{D}=\overline{W}_{\epsilon,p}^{p}$ (for some $\epsilon>0$). We denote by $(Y_{n},\pi_{n})$ an argmin in $A$ of
$(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right).$
Then we have
$\begin{split}&\mathbb{E}\left[\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}\right)-\underset{(Y,\pi)\in A}{\min}\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right]\\ &\leq C_{p,R}\sqrt{\frac{N(d+1)}{n}}+2K\left(\frac{e^{\frac{{\mathcal{K}}}{\epsilon}}}{\sqrt{n}}\left(1+\frac{1}{\epsilon^{\lfloor d/2\rfloor}}\right)\right),\end{split}$ (4.3)
with $\kappa=2C_{p}R+(2R)^{p}$, where $C_{p}=\sqrt{2}p(2R)^{p-1}$ is the Lipshitz constant of the Euclidean cost at the power $p$. Additionally, $K$ is an unknown constant depending only on $R$ and $p$ and $C_{p,R}$ is the constant associated to $W_{\epsilon,p}^{p}$ in Theorem 4.1.
As the regularization parameter $\epsilon$ goes to zero, the upper bound in (4.3) grows exponentially with the dimension. This arises from the estimation of $W_{\epsilon,p}(\mu^{l},\mu_{n}^{l})$. While the sample complexity of entropic optimal transport remains independent of the dimensionality (even for continuous measures), this no longer holds as $\epsilon\to 0$. In fact, in this limit the curse of dimensionality is implicitly hidden in the regularization parameter $\epsilon$.
@@ 5.1 Sample complexity of optimal transport divergences
Although in general empirical optimal transport suffers from the curse of dimensionality fournier2015rate, it has now been well established that the rate of convergence of the empirical optimal transport divergences between a measure $\mu$ and an empirical measure $\nu_{n}$ distributed over $n$ i.i.d random variables of common law $\nu$, depends on the intrinsic complexity of the measures $\nu$ and $\mu$. If at least one of them is discretely supported, the convergence rate is of order $O(1/\sqrt{n})$ sommerfeld2018inference; genevay2019sample; hundrieser2024empirical; nadjahi2020statistical; lin2021projection, which does not depend exponentially on the dimension of the underlying space $\mathbb{R}^{d}$. In other words, the sample complexity of optimal transport adapts to the measure with lowest complexity hundrieser2024empirical. Our main theorem allows to recover this $\frac{1}{\sqrt{n}}$ rate of convergence for empirical semi-discrete transport. By choosing $L=1$, we get that for any $\nu\in\mathcal{M}_{1}^{N}(\mathbb{B}_{R})$,
$\mathbb{E}|\mathbb{D}(\mu_{n},\nu)-\mathbb{D}(\mu,\nu)|\leq\mathbb{E}\left[\sup_{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}\left|\mathbb{D}\left(\mu_{n},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\mathbb{D}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\right],$
the upper bounds of Theorem 4.1 also hold for the empirical divergences themselves. We note however that they do not improve upon the literature.
@@ 5.2 Barycenters with sparse support
For a sparse barycenter with at most $N$ support points, and empirical target measures supported on $n$ random variables each, the generalization error bounds of our main Theorem 4.1 are of order $\sqrt{N/n}$, which is classical in semi-discrete optimal transport (see (del2024central, Theorem 2.10)). Interestingly, this rate is uniform in the number of measures $L$, in the regularization parameter $\epsilon$ for the Sinkhorn divergence case, and exhibits no curse of dimensionality (i.e no exponential dependency on $d$, the dimension of the ambient space of the target measures).
For a set of discrete target probability measures, a solution of the unconstrained Wasserstein barycenter problem (1.1) for $\mathbb{D}=W_{2}^{2}$ is also a discrete measure. Indeed, assuming that each measure $\mu^{\ell}$ is supported on $n^{\ell}$ points, and setting $N\geq\sum_{\ell=1}^{L}n^{\ell}-L+1$, problems (1.1) and (1.2) coincide: there exists an unconstrained Wasserstein barycenter which turns out to be sparse, supported on a set of maximal cardinality $\sum_{\ell=i}^{L}n^{\ell}-L+1$ (anderes2016discrete, Theorem 2). In this setting, heinemann2022randomized obtains the same rate of convergence $\sqrt{N/\min_{\ell}(n^{\ell})}$ for the unconstrained barycenter, which corresponds to the sparse barycenter considered in this work. This analysis is only limited to large values of $N$.
@@ 5.3 Tightness of the result
In this section, we investigate the optimality of the estimation error bound in Corollary 4.4 with respect to the sample size $n$ and the sparsity level $N$.
@@ 5.3.1 Lower bound of the estimation error
We aim to establish a lower bound for the estimation error, i.e., the left-hand side term in Corollary 4.4, in a worst-case sense over the choice of target measures. In bartlett2002minimax, the authors consider the quantization problem and a worst case lower bound of the following quantity
$\mathbb{E}\left[W_{2}^{2}\left(\mu,\sum_{i=1}^{N}\bar{\pi}_{i}\delta_{y_{i}^{n}}\right)-\underset{Y\in\mathbb{B}_{R}^{N}}{\min}W_{2}^{2}\left(\mu,\sum_{i=1}^{N}\bar{\pi}_{i}\delta_{y_{i}}\right)\right],$ (5.1)
where $(y_{1}^{n},\ldots,y_{N}^{n})$ denotes a minimizer of $(y_{1},\ldots,y_{N})\mapsto W_{2}^{2}(\mu_{n},\sum_{i=1}^{N}\bar{\pi}_{i}\delta_{y_{i}})$ and $\bar{\pi}$ minimizes the function $\pi\mapsto W_{2}^{2}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$. A direct extension of this result provides a lower bound for the sparse $2$-Wasserstein barycenter problem. The argument is an application of (bartlett2002minimax, Theorem 1), a stronger minimax statement which we only explicit for the empirical barycenter.
@@
For a fixed $N$, Lemma 5.1 implies that the barycenter functional at a sparse barycenter computed on a dataset of $nL$ observations cannot converge faster than $1/\sqrt{n}$ without additional assumptions. The proof of Lemma 5.1 can be found in Appendix D.2.
We observe the same dependency on $n$ of the estimation error for both the upper bound in (4.2) and the lower bound in (5.2). However the dependency in $N$ differs between these two results. The influence of $N$ on the estimation error depends in fact heavily on the target measures and their intrinsic dimension.

### v3

@@ 3 Sample complexity for sparse optimal transport barycenters
This section presents our main results : statistical generalization bounds for optimal transport barycenters. We draw inspiration from (gyorfi2002principles, Section 4.4), which treats the K-means functional (see Section 4.4). The following provides uniform generalization bounds for the empirical barycenter cost function, which most barycenter algorithms seek to minimize in practice (see Section 2.3).
@@
where the finite constant values $C_{p,R}$ and $C_{d,N}$ depend on the divergence $\mathbb{D}$ and are summarized in Table 1.
@@
We consider empirical samples of equal size $n$ for each measure, as well as uniform weights $\frac{1}{L}$ for simplicity. Our results hold for different sample sizes, replacing $n$ by the minimal sample size, and any weights $\lambda\in\Delta_{L}$.
@@
Complete details are provided in Appendix C. We treat here the case $\mathbb{D}=W_{p}^{p}$ for simplicity. Let $(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}$, the goal is to control, for each $\ell=1,\ldots,L$, the variation between empirical and population $W_{p}^{p}$ divergences. This is achieved by resorting to the dual formulation of $W_{p}^{p}$, in which the dual variables are vectors $w\in\mathbb{R}^{N}$. Lemma D.2 ensures that there exist functions $f_{Y}^{w}$ such that for each $\ell$ and any $\pi\in\Delta_{N}$
$\left|W_{p}^{p}\!\left(\mu_{n}^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-W_{p}^{p}\!\left(\mu^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\leq\sup_{w\in\mathbb{R}^{N}}\left|\mathbb{E}_{X\sim\mu_{n}^{\ell}}\!\left[f_{Y}^{w}(X)\right]-\mathbb{E}_{X\sim\mu^{\ell}}\!\left[f_{Y}^{w}(X)\right]\right|.$
Proposition B.1 provides a bound on the dual variables, $w\in[-K,K]^{N}$ for some constant $K$ independent of $Y$ and $\pi$. This allows us to invoke standard results from empirical process theory (e.g., (gine2021mathematical, Theorem 3.5.1)). ∎
@@
In the specific case where $\mathbb{D}$ is a distance, using the triangular inequality we have $|\mathbb{D}(\mu_{n}^{\ell},\nu)-\mathbb{D}(\mu^{\ell},\nu)|\leq\mathbb{D}(\mu_{n}^{\ell},\mu^{\ell})$. Statistical results for $\mathbb{D}$ fournier2015rate could then be used to derive error estimates. However, for $\mathbb{D}=W_{1}$ for example, the estimate is of order $n^{-1/d}$ (chewi2024statistical, Proposition 2.1), compared with $n^{-1/2}$ in Theorem 3.1.
@@
Let $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{B}_{R})$, for $R>0$. For some integer $n$, let $\mu_{n}^{1},\ldots,\mu_{n}^{L}$ be empirical measures supported over $n$ i.i.d random variables of respective law $\mu^{\ell}$, for all $\ell\in[\![1,L]\!]$. Let also $N\geq 1$ be some integer, $A$ a closed nonempty subset of $\mathbb{B}_{R}^{N}\times\Delta_{N}$, and let $\mathbb{D}=W_{p}^{p},\>W_{\epsilon,p}^{p}$ (for some $\epsilon>0$), $SW_{p}^{p}$ or $\text{max-}SW_{p}^{p}$. We denote by $(Y_{n},\pi_{n})$ an argmin in $A$ of the empirical barycenter cost $(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$ where $\operatorname{F}_{\mathbb{D}}$ is given in (2.5). Then we have :
$\mathbb{E}\left[\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}\right)-\underset{(Y,\pi)\in A}{\min}\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right]\leq 2C_{p,R}\sqrt{\frac{C_{d,N}}{n}},$ (3.1)
@@
Remark 3.5.
For $\mathbb{D}=W_{2}^{2}$ and $L=1$, the estimation error in Corollary 3.4 is upper bounded by $O(N/\sqrt{n})$ biau2008performance, independently of the dimension $d$, to be compared to our $O(\sqrt{dN/n})$ results.
@@
This leverages classical statistical learning theory arguments. By continuity of the functionals $(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}})$ and $(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}})$, both admit respective global minimizers in $A$, $(Y^{*},\pi^{*})$ and $(Y_{n}^{*},\pi_{n}^{*})$, with associated discrete measures $\nu^{*}:=\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}$ and $\nu_{n}^{*}:=\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}$. Then we have
@@
Theorem 3.1 and the sample complexity results of Sinkhorn divergences in genevay2019sample also yield an upper bound for the debiased Sinkhorn barycenters genevay2018learning; janati2020debiased, although with constants which have an exponential dependence in the dimension. The proof of the next corollary is also postponed in Appendix C.
@@
Let $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{B}_{R})$, for $R>0$. For some integer $n$, let $\mu_{n}^{1},\ldots,\mu_{n}^{L}$ be empirical measures supported over $n$ i.i.d random variables of respective law $\mu^{\ell}$, for all $\ell\in[\![1,L]\!]$. Let also $N\geq 1$ be an integer, $A$ a closed nonempty subset of $\mathbb{B}_{R}^{N}\times\Delta_{N}$ and for any $\epsilon>0$, set $\overline{W}_{\epsilon,p}^{p}(\mu,\nu)=W_{\epsilon,p}^{p}(\mu,\nu)-\frac{1}{2}(W_{\epsilon,p}^{p}(\mu,\mu)+W_{\epsilon,p}^{p}(\nu,\nu))$. We denote by $(Y_{n},\pi_{n})$ an argmin in $A$ of the empirical barycenter cost $(Y,\pi)\mapsto\operatorname{F}_{\overline{W}_{\epsilon,p}^{p}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$ where $\operatorname{F}_{\mathbb{D}}$ is given in (2.5). Then we have
$\begin{split}\mathbb{E}\left[\operatorname{F}_{\overline{W}_{\epsilon,p}^{p}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}\right)-\underset{(Y,\pi)\in A}{\min}\operatorname{F}_{\overline{W}_{\epsilon,p}^{p}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right]&\leq\frac{C_{\overline{W}}}{\sqrt{n}}\end{split}$ (3.2)
with $C_{\overline{W}}=C_{p,R}N(d+1)+2Ke^{\frac{{\mathcal{K}}}{\epsilon}}\left(1+\epsilon^{-\lfloor d/2\rfloor}\right)$ with $\mathcal{K}=2\sqrt{2}p(2R)^{p-1}R+(2R)^{p}$, $K$ is a constant depending only on $R$ and $p$, and $C_{p,R}$ is the constant related to $W_{\epsilon,p}^{p}$ in Theorem 3.1.
@@ 4.1 Sample complexity of optimal transport divergences
Empirical optimal transport suffers from the curse of dimensionality fournier2015rate, but it is now well established that its statistical convergence depends on the intrinsic complexity of the target measures hundrieser2024empirical. If one of them is discrete, the convergence rate is of actually of order $O(1/\sqrt{n})$ sommerfeld2018inference; genevay2019sample; hundrieser2024empirical; nadjahi2020statistical; lin2021projection. Theorem 3.1 allows to recover convergence rates for the quantity $\mathbb{E}|\mathbb{D}(\mu_{n},\nu)-\mathbb{D}(\mu,\nu)|$, for any $\nu\in\mathcal{M}_{1}^{N}(\mathbb{B}_{R})$, as well as for the empirical divergences themselves. These do not improve upon the literature.
@@ 4.2 Barycenters with sparse support
For a sparse barycenter with at most $N$ support points and sample size $n$, Theorem 3.1 provides generalization bounds of order $\sqrt{N/n}$, classical in semi-discrete optimal transport (del2024central, Theorem 2.10). Interestingly, this rate is uniform in the number of measures $L$, in the regularization parameter $\epsilon$ for the Sinkhorn divergence, without exponential dependency on the ambiant dimension $d$.
For $\mathbb{D}=W_{2}^{2}$, assuming that each measure $\mu^{\ell}$ is supported on $n^{\ell}$ points, and setting $N\geq\sum_{\ell=1}^{L}n^{\ell}-L+1$, problems (1.1) and (1.2) coincide: an unconstrained barycenter turns out to be sparse (anderes2016discrete, Theorem 2). In this setting, heinemann2022randomized obtains the same rate of convergence $\sqrt{N/\min_{\ell}(n^{\ell})}$ for the unconstrained barycenter, this analysis being limited to large $N$.
@@ 4.3 Tightness of the result
In this section, we discuss the optimality of the estimation error bound in Corollary 3.4 with respect to the sample size $n$ and the sparsity level $N$.
@@ 4.3.1 Lower bound of the estimation error
We aim to establish a lower bound for the estimation error, i.e., the left-hand side term in Corollary 3.4, in a worst-case sense over the choice of target measures. We rely on (bartlett2002minimax, Theorem 1) which provides minimax lower bounds for the optimal quantization cost ($L=1$ and $\mathbb{D}=W_{2}^{2}$). A direct extension of this result provides a lower bound for the sparse $2$-Wasserstein barycenter problem.
@@
The proof of Lemma 4.1 can be found in D.2. For a fixed NN, Lemma 4.1 implies that the $O(1/\sqrt{n})$ convergence rate for barycenter functional in 3.4 cannot be improved without further assumptions. However the dependency in NN differs.